Instructions: Solve each system by the substitution method.
What is the solution for 4x-y=-3
y=4x+3

Answer by checkley71(8403)

Y=4X+3
4X-(4X+3)=-3
4X-4X-3=-3
4X-4X=-3+3
0=0 SHOWS THAT THESE TWO EQUATIONS ARE IDENTICAL
PROOF
4X-Y=-3
-Y=-4X-3
Y=4X+3 &
Y=4X+3